OpenStudy (anonymous):

Tell if the events are independent or dependent. Explain your reasoning. A small box is filled with 50 marbles. There are 15 red marbles, 15 blue marbles, 10 purple marbles, 5 yellow marbles, and 5 black marbles in the box. A yellow marble is drawn from the box, is not replaced, and then a black marble is drawn.

OpenStudy (jacobbenvenutty):

I would say that this is an Independent set of events. The definition of an independent event is: When the occurrence of one event does not affect the likelihood that the other event will occur. The definition of an dependent event is: When the occurrence of one event does affect the likelihood that the other event will occur. Seeing as how a yellow marble and a black marble are drawn they do not effect eachother making it an independent event.

OpenStudy (jacobbenvenutty):

After reading the definitions of both what do you think it is?

ganeshie8 (ganeshie8):

Since we're NOT replacing the marble, it should be dependent right ?

OpenStudy (jacobbenvenutty):

I think the question was asking if drawing the black marble effected the yellow marble... Or a least thats how i read it

ganeshie8 (ganeshie8):

P(black) = 5/50 P(black | yellow) = 5/49

ganeshie8 (ganeshie8):

since P(black) != P(black | yellow), both events are dependent

ganeshie8 (ganeshie8):

?

OpenStudy (jacobbenvenutty):

oh... ok lol

OpenStudy (jacobbenvenutty):

i must have just not understood the question fully

ganeshie8 (ganeshie8):

yeah t'was bit twisted lol if the events are drawing black and drawing yellow without replacing then they look dependent...